OpenStudy (goformit100):

A censusman on duty visited a house which the lady inmates declined to reveal their individual ages, but said — “we do not mind giving you the sum of the ages of any two ladies you may choose”. Thereupon the censusman said — “In that case please give me the sum of the ages of every possible pair of you”. The gave the sums as follows : 30,33,41,58,66,69. The censusman took these figures and happily went away. How did he calculate the individual ages of the ladies from these figures.

OpenStudy (goformit100):

@TessaMarie08181997 @mikaela19900630 I personally tagged you too.

OpenStudy (anonymous):

The sum of all the ages is (30+33+41+58+66+69)/3=99 Looking which sum gives 99 gives 30+69+,33+66,41+58. the ages are \[a \le b \le c \le d \] this gives us a+b= 30 c+d=69 this solves to 11,19, 22,47

OpenStudy (goformit100):

Thankyou Madam

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

your welcome